OpenStudy (anonymous):

Preston read twenty-eight pages in 2 hours and Bethany read eighty-five pages in 5 hours. Which of the following statements is true? A)Preston read at a rate of one-fourteenth pages per hour. B)Bethany read at a rate of four hundred twenty-five pages per hour. C)Bethany read more than three times as fast as Preston. D)Preston read 3 pages per hour slower than Bethany.

OpenStudy (anonymous):

D is true

OpenStudy (anonymous):

just divide the number of pages read by how many hours so preston read 28/2=14 pages per hour and bethany read 85/5=17 pages per hour

OpenStudy (calculusxy):

The answer is D. pages divided by the hour 28/2=14 (Preston) 85/5=17 (Bethany) 17 (Bethany)- 14(Preston)= 3 pages slower per hour.